OpenStudy (blenchiki):

How do I solve this? Please and thank you. Find the value of x. If necessary, round to the nearest tenth

OpenStudy (blenchiki):

2x^2= 28-x

OpenStudy (sanoofy):

x= -4 2(4)^2=28-4 2(16)=28-(-4) 32=32 Therefore, LS=RS